# Interplay between altermagnetism and topological superconductivity in an unconventional superconducting platform

### v1

@@ III.2 Local density of states
$\rho_{i}(E)=\sum_{n\sigma}(\lvert u^{n}_{i\sigma}\rvert^{2}+\lvert v^{n}_{i\sigma}\rvert^{2})\delta(E-E_{n}),$ (8)
where $u^{n}_{i\sigma}$ and $v^{n}_{i\sigma}$ are the electronlike and holelike quasiparticle amplitudes, respectively. Here, $n$ and $\sigma$ are the eigenvalue and spin indices, respectively.
Figure 4: Panels (a) and (b) show local density of states (LDOS) at zero energy ($E=0$) of the helical $p$-wave superconductor coupled to the altermagnet (Eq. (1)) for the values of the altermagnet amplitude $J_{A}=0.3t$ and $J_{A}=0.65t$, respectively. Panels (c) and (d) display LDOS at $E=0$ of the chiral $p$-wave superconductor for the coupling constant $J_{A}=0.3t$ and $J_{A}=0.8t$, respectively. The inset in each panel shows the spectrum of the closest-to-zero-energy states in the corresponding phase. We here set $\mu=2t$ and $\Delta_{p}^{h}=\Delta_{p}^{c}=1.0$, and $t=1$.
Figure 5: Nearly flat band edge mode driven by altermagnet in the presence of a mixed helical and chiral superconductor. Panels (a) and (b) show ribbon spectrum for a mixed superconductor when the altermagnet amplitude is below ($J_{A}<J_{A}^{c}$) and above ($J_{A}>J_{A}^{c}$) the critical value ($J_{A}^{c}$) at which the band structure undergoes the transition. Panels (c) and (d) depict corresponding zero energy local density of states for $J_{A}<J_{A}^{c}$ and $J_{A}>J_{A}^{c}$, respectively. We set the parameters $\mu=2t$ and $\Delta_{p}^{h}=\Delta_{p}^{c}=t$, with $t=1$, in Eq. (1).
We first consider a helical $p$-wave superconductor. In Fig. 4, we calculate the LDOS $\rho_{i}(E)$ at zero energy ($E=0$) in the $xy$ plane by employing OBC along both the $x$ and $y$ directions, using the lattice Hamiltonian in Eq. (2) in real space. We then find close-to-zero-energy dispersive edge modes that are maximally localized along the two edges of the system for $J_{A}<J_{A}^{c}$, as shown in Fig. 4(a), with the plot of the eigenvalues in the inset of the same figure. In the helical case, altermagnetism drives the system from hosting dispersive Majorana modes in the regime of subcritical altermagnet amplitude to featuring MFEMs when $J_{A}>J_{A}^{c}$. Interestingly, these zero-energy MFEMs are again maximally localized only along the two edges of the system, shown in Fig. 4(b), with the closest-to-zero-energy modes shown in the inset.
@@
For a chiral $p$-wave superconductor, when $J_{A}<J_{A}^{c}$, as expected, dispersive Majorana edge modes are highly localized along the four edges of the system, as shown in Fig. 4(c). In contrast, in the case of strong altermagnet OP $J_{A}>J_{A}^{c}$, the edge modes disappear, i.e., they become completely delocalized, as shown in Fig. 4(d), and the bulk enters the nodal line SC phase, consistent with previous analyses of the band structure and the edge states.
@@ III.3 Mixed topological phase
We now analyze the topological edge states and the LDOS arising from the interplay between $d-$wave altermagnet and a mixed pairing superconductor with both chiral and helical $p$-wave components.
In Figs. 5(a) and 5(b), we show the ribbon spectrum calculated under the PBC along the $x$ direction and OBC along the $y$ direction. The spectra reveal dispersive edge modes in the regime of the subcritical altermagnet-superconductor coupling, $J_{A}<J_{A}^{c}$, which are characterized by both linear and quadratic dispersions, as illustrated in Fig. 5(a). Notably, for $J_{A}>J_{A}^{c}$, the system transitions to a mixed topological phase, which supports both linearly dispersing and nearly flat edge modes, as depicted in Fig. 5(b). This should be contrasted with the behavior of the edge states observed in the pure chiral and helical superconducting phases, as shown in Fig. 3.
The features of the mixed phase are further corroborated by the $E=0$ LDOS shown in Figs. 5(c) and 5(d), where the zero-energy state is highly localized along the edges of the finite system, with the closest-to-zero-energy states depicted in the inset. This suggests that the $d-$wave symmetry of altermagnetism facilitates the formation of this mixed topological phase, which accommodates both dispersive and flat edge states in the presence of mixed helical and chiral $p-$wave pairings. Such behavior is clearly distinct from that observed in either the purely chiral or helical phase, as can be observed in Fig. 4.
@@
In this section, we analyze the variation of the bulk gap profile by employing PBC along both the $x$ and $y$ directions as a characteristic signature of the topological phase transition.
As an indicator of the topological phase transition, in Fig. 6, we show the maximum value of the gap, $\Delta=|E_{2}-E_{1}|$, between closest to zero energy Bogoliubov quasiparticle bands, where $E_{1}$ ($E_{2}$) represents the maximum (minimum) of the energy of the negative (positive) energy band. Fig. 6(a) displays the gap, $\Delta$, as a function of the altermagnet strength $J_{A}$ and chemical potential $\mu$ for the helical $p$-wave superconductor. For a fixed (nonzero) value of $\mu$, we observe that this gapped topologically non-trivial phase remains robust in the regime $J_{A}<J_{A}^{c}$, where $J_{A}^{c}$ is the critical coupling for the phase transition at the zero chemical potential. On the other hand, for a strong altermagnet amplitude, $J_{A}>J_{A}^{c}$, we observe multiple topological phase transitions in terms of the chemical potential, with the bulk gap closing and reopening. Each gap closing point corresponds to a gapless topological superconducting phase hosting MFEMs, as depicted in Fig. 3(b).
In Fig. 6(b), we display the outcome of an analogous analysis for a chiral $p$-wave superconductor. For a fixed value of $\mu$, we observe a topological phase transition from a topological chiral gapped superconducting phase to a gapless nodal-line superconducting phase, see also Fig. 3(d). Finally, in Fig. 6(c), we show the maximum gap $\Delta$ in the $\mu-J_{A}$ plane for a mixed $p$-wave superconductor. Interestingly, in this case, we also find multiple topological phase transitions between gapped and gapless phases, which is a clear indication of a mixed topological phase that hosts MFEMs and dispersive edge states, shown in Fig. 5(b).
@@ V Summary and discussion
Our findings, particularly, the predicted gapless topological phase and MFEMs may be relevant in light of a recent experimental proposal [31], where a gapless topological superconducting phase hosting MFEMs in a magnet-superconductor heterostructure composed of Fe/Ta(110) has been identified. Here, we propose an alternative method to engineer gapless topological superconductivity hosting MFEMs in an altermagnet and unconventional superconductor hybrid system. Specifically, the potential experimental setup realizing our scenario involves $\rm{Ru}\rm{O}_{2}$, a leading candidate for altermagnetism with $d$-wave magnetic order [1], fabricated on top of a topological $p$-wave superconductor, such as $\rm{U}\rm{Te}_{2}$ [48]. This heterostructure could serve as a promising candidate for engineering gapless topological superconductivity hosting MFEMs. Although this remains experimentally challenging due to the limited availability of topological $p-$wave superconducting materials, our study opens new avenues for engineering gapless, and, particularly, hybrid topological SC states, wherein Majorana dispersive and flat edge modes coexist within an unconventional SC platform with zero net magnetization.
@@ Acknowledgments
PC acknowledges Arijit Saha, Kenji Fukushima, and Takashi Oka for fruitful discussions and support. V. J. acknowledges the support of the Swedish Research Council Grant No. VR 2019-04735 and Fondecyt (Chile) Grant No. 1230933.

### v2

@@ III.2 Local density of states
$\rho_{i}(E)=\sum_{n\sigma}\left(\lvert u^{n}_{i\sigma}\rvert^{2}+\lvert v^{n}_{i\sigma}\rvert^{2}\right)\delta(E-E_{n}),$ (9)
where $u^{n}_{i\sigma}$ and $v^{n}_{i\sigma}$ are the electron-like and hole-like quasiparticle amplitudes, respectively, and $n$ and $\sigma$ denote the eigenstate and spin indices.
We first consider a helical $p$-wave superconductor. In Fig. 4, we compute the zero-energy LDOS $\rho_{i}(E=0)$ across the two-dimensional plane by implementing open boundary conditions (OBC) along both the $x$ and $y$ directions and diagonalizing the lattice Hamiltonian in Eq. (2). For subcritical values of the altermagnetic coupling, $J_{A}<J_{A}^{c}$, the system hosts dispersive edge modes that are sharply localized along the two boundaries along the $y$-direction, with a negligible LDOS along the $x$-edge, as shown in Fig. 4(a), which is also consistent with the low-energy edge theory, as discussed in App. B. The corresponding eigenvalue spectrum, displayed in the inset, confirms the presence of Majorana zero-energy modes. These dispersive Majorana edge states are a hallmark of the helical topological phase, which remains protected by the particle-hole symmetry (PHS) in the presence of weak altermagnetic perturbations.
As the altermagnetic strength is increased beyond the critical value $J_{A}^{c}$, the system undergoes a topological phase transition into a gapless regime featuring MFEMs. These MFEMs manifest as the states close to zero energy featuring weak dispersion, as illustrated in Fig. 4(b). Remarkably, despite the loss of TRS due to the altermagnet, the flat modes remain maximally localized along the same edges, indicating a robust topological character that survives the gap-closing transition. The flatness of these modes in energy signals the collapse of the bulk gap at isolated points in momentum space and the emergence of a weak topological phase composed of symmetry-protected one-dimensional chains.
In the chiral $p$-wave phase, the LDOS again reveals a distinct topological character. For $J_{A}<J_{A}^{c}$, we observe the expected behavior: dispersive chiral Majorana modes are strongly localized along all four edges of the system, consistent with the nonzero Chern number of the chiral SC phase [Fig. 4(c)]. For $J_{A}>J_{A}^{c}$, the LDOS profile changes drastically, as shown in Fig. 4(d), where we can observe that the edge localization is lost, and the zero-energy spectral weight is shifted to the bulk. This feature further corroborates the onset of a nodal-line SC phase, with its signatures previously found in the band structure and edge state analyses.
Figure 5: Nearly flat band edge mode driven by altermagnet in the presence of a mixed helical and chiral superconductor. Panels (a) and (b) show ribbon spectrum for a mixed superconductor with the amplitude of the altermagnet OP $J_{A}=0.3t$ ($J_{A}<J_{A}^{c}$) and $J_{A}=0.65t$ ($J_{A}>J_{A}^{c}$) respectively, where the critical value is $J_{A}^{c}=0.5t$ at which the band structure undergoes the transition. Panels (c) and (d) depict corresponding zero energy local density of states for $J_{A}<J_{A}^{c}$ and $J_{A}>J_{A}^{c}$, respectively. The inset in panels (c) and (d) shows the spectrum of the closest-to-zero-energy states in the corresponding phase. We set the parameters $\mu=2t$ and $\Delta_{p}^{h}=\Delta_{p}^{c}=t$, with $t=1$, in Eq. (1).
@@ III.3 Mixed topological phase
We now analyze the topological edge states and the LDOS arising from the interplay between $d$-wave altermagnetism and a mixed pairing superconductor with both chiral and helical $p$-wave components.
In Figs. 5(a) and 5(b), we show the ribbon spectrum calculated under periodic boundary conditions (PBC) along the $y$ direction and open boundary conditions (OBC) along the $x$ direction. The spectra reveal rich edge-state structures in both the subcritical ($J_{A}<J_{A}^{c}$) and overcritical $J_{A}>J_{A}^{c}$ regimes of the altermagnetic coupling. In the former regime, the system supports multiple dispersive edge modes, which exhibit both linear and quadratic dispersions, as shown in Fig. 5(a). These modes emerge as a consequence of the simultaneous presence of TRS-preserving (helical) and TRS-breaking (chiral) pairing components, which can hybridize due to the breaking of this symmetry. In the mixed case, both chiral and helical edge states exhibit linear dispersion close to zero edge momentum for $J_{A}=0$. Upon coupling to the altermagnet, quadratically dispersing modes may emerge alongside the linear ones due to the quadratic momentum dependence of the altermagnet coupling near zero momentum, given its form in Eq. (4). On the other hand, the linearly dispersing modes in the mixed phase are due to the fully spin polarized form of the chiral SC pairing ($\sim s_{0}$). See Fig.5(a).
Notably, when the altermagnet strength exceeds the critical coupling, $J_{A}>J_{A}^{c}$, the edge spectrum undergoes a qualitative change. As seen in Fig. 5(b), the system enters a mixed topological phase characterized by the coexistence of both linearly dispersing and nearly flat edge modes, which can be qualitatively explained as follows. The quadratically dispersing mode in the weak-coupling regime due to an enhanced $E=0$ DOS and the lack of topological protection is rendered unstable at strong coupling. Since the chosen $d$-wave altermagnetic OP does not open a full gap at the edges [the $y$-edge remains gapless, as implied by Eq. (14), and explicitly shown in App. (B)], to minimize the kinetic energy , the spectral density of the quadratic band is transferred to low energies while forming an almost flat band. Such nearly flat modes feature localized, zero-energy MFEMs. On the other hand, the linearly dispersive branches at the center of the Brillouin zone remain gapless at strong coupling due to their chiral nature and the fact that the chosen altermagnetic coupling does not open an edge gap in the chiral case.
The presence of both flat and dispersive edge states in this mixed regime is further supported by the $E=0$ LDOS profiles, shown in Figs. 5(c) and 5(d). For both subcritical and supercritical values of $J_{A}$, the zero-energy states are localized only at the $y$-edge of the system. The inset plots display the low-energy eigenvalue spectrum, confirming the presence of quasi-zero-energy states.
@@
To gain further insight in the phase diagram of the model, in this section, we analyze the variation of the bulk gap profile by employing PBC along both the $x$ and $y$ directions as a characteristic signature of the topological phase transition.
We calculate the topological invariant for the purely chiral and helical phases in the Hamiltonian (2) for the weakly coupled regime $J_{A}<J_{A}^{c}$. The Chern number can be expressed as [47, 48]
$C=\frac{1}{2\pi i}\smashoperator[]{\int_{\rm BZ}^{}}\!d^{2}k\,\text{Tr}\Big{(}P^{-}\big{[}\partial_{k_{x}}P^{-},\partial_{k_{y}}P^{-}\big{]}\Big{)}$ (10)
in terms of the projector to the two filled bands $P^{-}=P_{1}^{-}+P_{2}^{-}$, where
$\begin{split}P_{1}^{-}=\frac{(E_{1}-H)(E_{2}-H)(E_{-2}-H)}{(E_{1}-E_{-1})(E_{2}-E_{-1})(E_{-2}-E_{-1})},\\ P_{2}^{-}=\frac{(E_{1}-H)(E_{2}-H)(E_{-1}-H)}{(E_{1}-E_{-2})(E_{2}-E_{-2})(E_{-1}-E_{-2})}.\end{split}$ (11)
Here, $E_{-1}$ and $E_{-2}$ are the energies of the two filled bands, and $E_{1}$ and $E_{2}$ are the energies of the unoccupied bands. In both cases, the Chern number turns out to be equal to one, therefore corroborating the topological nature of these phases. To distinguish between the gapped and gapless topological SC phases, we calculate the bandgap and analyze the form of the edge states.
As an indicator of the topological phase transition, in Fig. 6, we show the maximum value of the gap, $\Delta=|E_{2}-E_{1}|$, between closest to zero energy Bogoliubov quasiparticle bands, where $E_{1}$ ($E_{2}$) represents the maximum (minimum) of the energy of the negative (positive) energy band. Fig. 6(a) displays the gap, $\Delta$, as a function of the altermagnet strength $J_{A}$ and chemical potential $\mu$ for the helical $p$-wave superconductor. For a fixed (nonzero) value of $\mu$, we observe that this gapped topologically non-trivial phase remains robust in the regime $J_{A}<J_{A}^{c}$, where $J_{A}^{c}$ is the critical coupling for the phase transition at the zero chemical potential. On the other hand, for a strong altermagnet amplitude, $J_{A}>J_{A}^{c}$, we observe multiple topological phase transitions in terms of the chemical potential, with the bulk gap closing and reopening. Each gap closing point corresponds to a gapless topological SC phase hosting MFEMs, as depicted in Fig. 3(b).
In Fig. 6(b), we display the outcome of an analogous analysis for a chiral $p$-wave superconductor. For a fixed value of $\mu$, we observe a topological phase transition from a chiral gapped SC phase to a gapless nodal-line SC phase, see also Fig. 3(d). Finally, in Fig. 6(c), we show the maximum gap $\Delta$ in the $\mu-J_{A}$ plane for a mixed $p$-wave superconductor. Interestingly, in this case, we also find multiple topological phase transitions between gapped and gapless phases, which is a clear indication of a mixed topological phase that hosts MFEMs and dispersive edge states, shown in Fig. 5(b).
@@ V Summary and discussion
In this work, our primary objective was to identify signatures of flat and mixed (dispersive + flat) edge modes in an altermagnet-superconductor hybrid system. Based on the $d$-wave nature of the altermagnet OP, we expect that the edge states can be controlled by orientation of the altermagnet, characterized by the azimuthal and polar angles, $\theta$ and $\phi$, respectively. To this end, in our model, the altermagnet couples to spin along the $x$-direction, i.e., $\theta=\pi/2$, $\phi=0$, which results in a coupling term that neither commutes nor anticommutes with the Hamiltonian of parent $p$-wave topological superconductor. This specific form leads to selective gapping of the corresponding edge-state Hamiltonian. Consequently, the edge states are localized along the $y$-edges, corresponding to $L_{x}=1$ and $L_{x}=30$, in the mixed (helical case), as shown in Figs. 5(c) and (d) [4(a) and (b)]. Notably, because the superconductor-altermagnet coupling along the $x$-direction does not constitute a mass term for the parent Hamiltonian, the emergence of higher-order topological phases is precluded in our setup. This should be contrasted with the situation in Ref. [14], where a different form of the coupling enables a sign-changing Dirac mass, thereby allowing for the realization of a higher-order topological SC phase in a parent helical $p$-wave superconductor.
Our findings, particularly, the predicted gapless topological phase and MFEMs may be relevant in light of a recent experimental proposal [35], where a gapless topological SC phase hosting MFEMs in a magnet-superconductor heterostructure composed of Fe/Ta(110) has been identified. Here, we propose an alternative method to engineer gapless topological superconductivity hosting MFEMs in such a hybrid system. Specifically, the potential experimental setup realizing our scenario involves $\rm{Ru}\rm{O}_{2}$, a leading candidate for altermagnetism with $d$-wave magnetic order [1] (see, however, Ref. [49]), fabricated on top of a topological $p$-wave superconductor, such as $\rm{U}\rm{Te}_{2}$ [50]. This heterostructure could serve as a promising candidate for engineering gapless topological superconductivity hosting MFEMs. Although this remains experimentally challenging due to the limited availability of topological $p$-wave SC materials, our study opens new avenues for engineering gapless, and, particularly, hybrid topological SC states, wherein Majorana dispersive and flat edge modes coexist within an unconventional SC platform featuring zero net magnetization. Finally, our work should motivate further studies on the characterization of the transition between gapped and nodal-point or nodal-line superconductor, in particular the topological and thermodynamic features of such transitions [51, 52].
@@ Acknowledgments
PC acknowledges Arijit Saha, Kenji Fukushima, and Takashi Oka for fruitful discussions and support. V. J. acknowledges the support of Fondecyt (Chile) Grant No. 1230933 and the Swedish Research Council Grant No. VR 2019-04735.
Data Availability
The data that support the findings of this study are available from the authors upon reasonable request.
Appendix A Symmetries of the Bogoliubov-de Gennes Hamiltonian in Eq. (2)
In this Appendix, we consider the discrete symmetries of the BdG Hamiltonian introduced in Eq. (2), anti-unitary particle-hole symmetry (PHS), time-reversal symmetry (TRS), and chiral symmetry (CS) (or unitary PHS). The results of this analysis are summarized in Table 1.
A.1 Anti-unitary particle-hole symmetry
The BdG Hamiltonian features PHS by construction, with the corresponding anti-unitary operator $\mathcal{P}$ acting on the Nambu spinor as
$\mathcal{P}=\tau_{x}s_{z}\mathcal{K},$
where $\mathcal{K}$ denotes complex conjugation. The symmetry condition for the Hamiltonian $h_{\bf k}$ reads as
$\mathcal{P}h_{\bf k}\mathcal{P}^{-1}=-h_{-{\bf k}}.$
Using the structure of $h_{\bf k}$ in Eq. (2), we find that terms transform under $\mathcal{P}$ as follows:
  1. (a)
    $\xi_{\bf k}\tau_{z}s_{0}\rightarrow-\xi_{-{\bf k}}\tau_{z}s_{0}$ (since $\xi_{\bf k}=\xi_{-{\bf k}}$),
  2. (b)
    $J({\bf k})\tau_{0}s_{x}\rightarrow-J({\bf k})\tau_{0}s_{x}$ (with $J({\bf k})=J(-{\bf k})$),
  3. (c)
    pairing terms $\sim\sin k_{x,y}$ are odd in momentum multiplied by the matrices that commute with the particle-hole operator, $\mathcal{P}=\tau_{x}s_{z}\mathcal{K}$, for both helical and chiral SC OPs.
Hence, the Hamiltonian in Eq. (2) possesses anti-unitary PHS with $\mathcal{P}^{2}=+1$.
A.2 Time-Reversal Symmetry (TRS).
The time-reversal operator for spin-$1/2$ systems is given by
$\mathcal{T}=\tau_{0}(is_{y})\mathcal{K},$
with the TRS condition:
$\mathcal{T}h_{\bf k}\mathcal{T}^{-1}=h_{-{\bf k}}.$
(a) Chiral $p$-wave case ($\Delta_{p}^{h}=0$):
The chiral pairing term
$\Delta_{p}^{c}(\tau_{x}s_{x}\sin k_{x}-\tau_{y}s_{x}\sin k_{y})$
is odd under time reversal:
$\mathcal{T}\tau_{x}s_{x}\mathcal{T}^{-1}$ $=-\tau_{x}s_{x},$
$\mathcal{T}\tau_{y}s_{x}\mathcal{T}^{-1}$ $=-\tau_{y}s_{x},$
$\sin k_{x,y}$ $\rightarrow-\sin k_{x,y}.$
The altermagnetic term $J({\bf k})\tau_{0}s_{x}$ also breaks TRS:
$\mathcal{T}\tau_{0}s_{x}\mathcal{T}^{-1}=-\tau_{0}s_{x},\quad J({\bf k})=J(-{\bf k}).$
Therefore, the chiral $p$-wave superconductor explicitly breaks TRS, even without the altermagnetic term.
(b) Helical $p$-wave case ($\Delta_{p}^{c}=0$):
The helical pairing term is
$\Delta_{p}^{h}(\tau_{y}s_{0}\sin k_{x}-\tau_{x}s_{z}\sin k_{y}).$
Under TRS:
$\mathcal{T}\tau_{y}s_{0}\mathcal{T}^{-1}$ $=-\tau_{y}s_{0},$
$\mathcal{T}\tau_{x}s_{z}\mathcal{T}^{-1}$ $=\tau_{x}s_{z},$
$\sin k_{x,y}$ $\rightarrow-\sin k_{x,y}.$
Thus, both terms in the helical pairing are even under TRS. However, the altermagnetic term $J({\bf k})\tau_{0}s_{x}$ again breaks TRS, as in the chiral case. Therefore, while the helical superconductor alone preserves TRS, the inclusion of altermagnetism leads to TRS breaking.
A.3 Chiral Symmetry (CS).
Chiral or unitary particle-hole symmetry is defined as the combination of anti-unitary PHS and TRS:
$\mathcal{C}=\mathcal{P}\mathcal{T}=\tau_{x}s_{z}\cdot\tau_{0}is_{y}=\tau_{x}s_{x}.$
The symmetry condition is:
$\mathcal{C}h_{\bf k}\mathcal{C}^{-1}=-h_{\bf k}.$
However, since the Hamiltonian with altermagnetism breaks TRS, chiral symmetry is absent in both the helical and chiral cases.
In summary, the BdG Hamiltonians with altermagnetic order, supporting both chiral and helical SC OPs, belong to symmetry class D in the tenfold classification of topological insulators and superconductors. This class supports a $\mathbb{Z}$ topological classification in two spatial dimensions [26] and therefore implies the possible existence of topologically protected chiral Majorana edge modes.
Appendix B Edge-state analysis: Purely helical pairing
To provide an analytic understanding of the edge-state structure in the presence of helical $p$-wave SC OP, we consider the low-energy expansion of the Bogoliubov–de Gennes (BdG) Hamiltonian in Eq. (2) with the chiral pairing $\Delta_{p}^{c}=0$. We analyze the edge states by considering a continuum limit near the $X$-point in the Brillouin zone at the momentum, $\mathbf{K}_{X}=(\pi,0)$, and imposing semi-infinite boundary conditions, therefore considering: (1) the $y$-edge, with the $x$-direction open and the $k_{y}$ momentum conserved; (2) the $x$-edge, with the $y$-direction open and the $k_{x}$ momentum conserved.
To this end, we first expand the Hamiltonian (2) about the $X$-point in the Brillouin zone to obtain,
$h_{X}(\mathbf{k})$ $=\left[-t(k_{x}^{2}-k_{y}^{2})+4t-\mu\right]\tau_{z}s_{0}$
$+[-4J_{A}+J_{A}(k_{x}^{2}+k_{y}^{2})]\tau_{0}s_{x}-\Delta_{p}^{h}\left(k_{x}\tau_{y}s_{0}+k_{y}\tau_{x}s_{z}\right).$ (12)
Here, the $\mathbf{k}$ is a small deviation of the momentum from the $X$-point, $|\mathbf{k}|\ll|\mathbf{K}_{X}|$.
B.1 Edge parallel to the $y$-direction
We now replace $k_{x}\to-i\partial_{x}$, keeping $k_{y}$ as a good quantum number in Eq. (B)
$H_{y\text{-edge}}$ $=\left[t(k_{y}^{2}+\partial_{x}^{2})+4t-\mu\right]\tau_{z}s_{0}+J_{A}(-4-\partial_{x}^{2}+k_{y}^{2})\tau_{0}s_{x}$
$+\Delta_{p}^{h}\left(i\partial_{x}\tau_{y}s_{0}-k_{y}\tau_{x}s_{z}\right).$ (13)
We look for the zero energy solutions of the form $\psi(x)=e^{-\kappa x}\chi$, such that ${\rm Re}(\kappa)>0$ ensuring normalizability of the mode, which subject to open boundary conditions at $x=0$, $\psi(x=0)=0$. To find such a zero mode we take $k_{y}=0$ and the low-energy effective edge Hamiltonian then takes reduces to
$H_{y\text{-edge}}$ $=\left[t\kappa^{2}+4t-\mu\right]\tau_{z}s_{0}-J_{A}(4+\kappa^{2})\tau_{0}s_{x}+i\Delta_{p}^{h}\kappa\tau_{y}s_{0}.$ (14)
The inverse localization length of the zero energy Majorana modes, $\kappa$, is obtained from the condition of the vanishing determinant of this matrix. We then find
$\mathrm{det}\,H_{y\text{-edge}}$ $=(J_{A}^{2}-t^{2})\kappa^{4}+(8J_{A}^{2}-8t^{2}+\Delta_{h}^{2}+2t\mu)\kappa^{2}$
$\quad+(16J_{A}^{2}-16t^{2}+8t\mu-\mu^{2})$ (15)
which we will not analyze in detail. We just point out that for the values of the parameters in Figs, 3(a) and 4(a) we find a pair of localized zero modes, which indeed yield a pair of counterpropagating edge modes obtained by the diagonalization of the Hamiltonian (2) in a semi-infinite geometry, with the spectrum shown in Fig. 3(a). This behavior is also expected based on the form of the edge Hamiltonian (14) where the altermagnetic term commutes with the pairing term, which is the effective kinetic term in the edge Dirac Hamiltonian. Therefore, such a term is expected not to gap them out, and these modes being topological should remain at zero energy.
B.2 Edge parallel to the $x$-direction
Next, we consider a $x$-edge geometry and perform a similar expansion:
$H_{x\text{-edge}}$ $=\left[-t(k_{x}^{2}+\partial_{y}^{2})+4t-\mu\right]\tau_{z}s_{0}+J_{A}(-4-\partial_{y}^{2}+k_{x}^{2})\tau_{0}s_{x}$
$\quad-\Delta_{p}^{h}\left(k_{x}\tau_{y}s_{0}-i\tau_{x}s_{z}\partial_{y}\right).$ (16)
Using the ansatz $\psi(y)=e^{-\kappa y}\chi$, the projected effective Hamiltonian for the $x$-edge for the zero mode becomes
$H_{x\text{-edge}}=\left(-t\kappa^{2}+4t-\mu\right)\,\tau_{z}s_{0}+\left(-4J_{A}-J_{A}\kappa^{2}\right)\,\tau_{0}s_{x}$
$-i\,\Delta_{h}\kappa\,\tau_{x}s_{z},$ (17)
which is obtained after setting $k_{x}=0$ in Eq. (B.2). The corresponding condition for the zero modes is obtained by setting the determinant of this Hamiltonian to zero, which explicitly reads
$\ {\det}(H_{x\text{-edge}})=J_{A}^{4}(4+\kappa^{2})^{4}$
$\ +\left[t^{2}(-4+\kappa^{2})^{2}+2t(-4+\kappa^{2})\mu+\mu^{2}-\Delta_{h}^{2}\kappa^{2}\right]^{2}$
$\ -2J_{A}^{2}(4+\kappa^{2})^{2}\left[\Delta_{h}^{2}\kappa^{2}+t^{2}(-4+\kappa^{2})^{2}+2t(-4+\kappa^{2})\mu+\mu^{2}\right]$ (18)
For the regime of parameters in Figs. 3(a) and 4(a), the above expression yields four distinct positive purely real values for the inverse localization length $\kappa$ (besides four negative ones, which correspond to non-normalizable modes). However, these modes are incompatible with the open boundary condition at $y=0$, and therefore no zero modes exist in this case. This outcome may be expected also based on the fact that the altermagnetic term anticommutes with the pairing term in the effective edge Hamiltonian (B.2), and therefore it is expected to gap out the edge modes.
These results are in agreement with our numerical findings displayed in Fig. 3(a) and 4(a).
The case of the edge states in the chiral SC can be analyzed in a similar manner but we do not delve into this case here.